OpenStudy (anonymous):

an object movws along the x-axis with its position x, in meters, given as a function of time t, in seconds, by x(t)=1.39t^2-9.25t+4.68. what is the object's velocity at time t=1.55s ?

OpenStudy (abhisar):

Hello @avemariab93 ! Welcome to OpenStudy !

OpenStudy (abhisar):

\(\sf Velocity = \frac{\huge dx}{\huge dt}\) => Velocity = \(\sf \frac{\huge d(1.39t^2-9.25t+4.68)}{\huge dt}=2.78t-9.25\). Now substitute the value of t=1.55 in the equation 2.78t-9.25.

OpenStudy (anonymous):

@abhisar omg!! thanks so much!! you're a life saver ^^

OpenStudy (abhisar):

Do yo know ? You can reward those who help you by hitting the blue best response button. You're most welcome :)